# NewtonBench: Benchmarking Generalizable Scientific Law Discovery in LLM Agents

### v1

@@ 4.1 Can LLM Agents Perform Generalizable Scientific Law Discovery?
Our benchmark proves highly challenging for models with weaker reasoning abilities. All three non-thinking LLMs (GPT-4.1-mini, GPT-4.1, and DeepSeek-V3) achieve overall symbolic accuracies below 10%. While these models show limited proficiency (20-50% accuracy) even in the simplest setting, their performance degrades precipitously in all other configurations, demonstrating that success in NewtonBench is contingent on strong, generalizable reasoning.
@@
Most reasoning models achieve overall accuracies between 30-70%, with GPT-5 and Gemini-2.5-pro distinguishing themselves at 72.9% and 65.0%, respectively. While these models all demonstrate proficiency in the easiest settings (80-100% accuracy), a substantial performance gap emerges as the difficulty escalates. In the most challenging setting, GPT-5 and Gemini-2.5-pro maintain accuracies of 29.9% and 13.9%, whereas all other reasoning models fall below 5%. This stark performance gap demonstrates that the benchmark’s difficulty controls are highly effective at stress-testing and stratifying advanced reasoning abilities.
@@ 4.2 Sensitivity to Observational Noise
To investigate robustness against noisy observations, we conducted an experiment on GPT-5-mini with four levels of Gaussian noise (0.0001, 0.001, 0.01, and 0.1). As illustrated in Figure 3, performance degrades significantly even with minimal noise, with the introduction of just a 0.0001 noise level causing a 13-15% reduction in accuracy compared to the ideal, noise-free setting. As the noise level was increased from 0.0001 to 0.1, symbolic accuracy declined proportionally, while data fidelity (RMSLE) remained relatively stable. Notably, code assistance did not affect noise robustness, with the baseline and code-assisted agents degrading similarly. This demonstrates that the symbolic accuracy of LLMs is extremely fragile to noise in observational data.
@@ 4.3 Cross-Domain Performance Disparities
Performance across the physics domains, illustrated in Figure 3, reveals a substantial difference in average accuracy, with scores ranging from 18% to 54%. Bose-Einstein Distribution, the most advanced and obscure domain in our benchmark, yields the lowest average accuracy at 18.1%. Furthermore, the performance disparities are exacerbated as system complexity increases. For instance, in the simple setting, Fourier’s Law yields 58.1% accuracy, comparable to the easiest domain, Acoustic Velocity (60.4%). In the complex setting, however, accuracy for Fourier’s Law plummets to 2.3%, while Acoustic Velocity remains at 45.0%. This demonstrates that the intrinsic nature of a physical domain is a primary factor governing the difficulty of model discovery. Full domain analysis is provided in Appendix E.3.
@@ 4.5 The Paradox of Code Assistance: The Exploration-Exploitation Trade-Off
To investigate this hypothesis, we adapt a common analysis approach for reasoning LLMs (Wang et al., 2025b; a) by identifying signature tokens associated with exploration (e.g., What if, Alternatively) and exploitation (e.g., Confirm, Refine). From these, we calculate the exploration rate (the percentage of exploration tokens among all such planning tokens; see Appendix E.4 for details). Figure 6(b) shows the results for two reasoning LLMs where reasoning traces are accessible. We observe a sharp drop in the exploration rate for Gemini-2.5-flash as the code budget increases from zero to one, while the rate remains stable for Qwen-3-235b. This suggests the performance degradation in stronger models reflects over-exploitation [^4] These observations are correlational and should be interpreted as suggestive rather than causal..
@@ 5 Related Work
Symbolic regression (SR) aims to discover mathematical formulas from data, with foundational approaches being Genetic Programming (GP) that evolve populations of candidate expressions (Koza, 1994; Augusto & Barbosa, 2000; Billard & Diday, 2002; Schmidt & Lipson, 2009). Meanwhile, more advanced approaches have been proposed, from physics-inspired, Pareto-driven strategies (Udrescu & Tegmark, 2020; Udrescu et al., 2020) to deep learning-based models (Petersen et al., 2021; Kamienny et al., 2022). The latest evolution in SR leverages the reasoning capabilities of LLMs to hypothesize and refine scientific equations (Shojaee et al., 2025a; b).
@@
The agentic capabilities of LLMs are increasingly being applied to automate the scientific process (Zheng et al., 2025; Wei et al., 2025), from assisting with experimental procedures in chemistry and biomedicine (Gottweis et al., 2025; Yang et al., 2025; Luo et al., 2025; Chen et al., 2025) to automating complex research workflows in machine learning (Chan et al., 2025; Huang et al., 2024a; Jiang et al., 2025; Jansen et al., 2025). This trend culminates in the pursuit of a fully autonomous “AI Scientist” capable of managing the entire research pipeline for open-ended discovery (Lu et al., 2024; Yamada et al., 2025).
@@
Virtual environments for LLM agents have evolved from general-purpose sandboxes for tasks like instruction following (Shridhar et al., 2021), web navigation (Yao et al., 2023), and compositional planning (Prasad et al., 2024), to specialized platforms for scientific reasoning, ranging from curriculum-based tasks (Wang et al., 2022) to open-ended discovery (Jansen et al., 2024).
@@ 6 Conclusion
NewtonBench is designed exclusively for benchmarking and evaluating the scientific discovery capabilities of LLM agents in a fully contamination-free, virtual environment. The metaphysical shifts applied to physical laws ensure the tasks are novel and not directly suitable for training LLMs to discover real-world physics. Due to the interactive, model‑system nature of the tasks, we exclude traditional symbolic regression methods that do not support model systems and LLM‑based symbolic regression pipelines whose prior reliance and workflows are incompatible with our protocol; these are outside the scope of this evaluation. We caution that the benchmark is not intended as a training dataset for developing stronger physics discovery models, but rather as a controlled tool for fair and rigorous evaluation. All shifted laws were curated by expert co‑authors (Physics Olympiad backgrounds) and cross‑validated.
@@ A.1 LLM Details
  • GPT-4.1-mini (OpenAI, 2025a) is a lightweight proprietary non-reasoning LLM from OpenAI, designed for efficient inference.
  • GPT-4.1 (OpenAI, 2025a) is the latest flagship proprietary non-reasoning LLM from OpenAI.
  • o4-mini (OpenAI, 2025c) is a proprietary reasoning LLM developed by OpenAI, optimized for lightweight inference and efficient deployment.
  • GPT-5-mini (OpenAI, 2025b) is a lightweight reasoning LLM from OpenAI, aiming for improved efficiency on downstream tasks.
  • GPT-5 (OpenAI, 2025b) is the strongest reasoning LLM by OpenAI with advanced reasoning capabilities.
  • DeepSeek-V3 (DeepSeek-AI, 2024) is an open-source non-reasoning LLM released by DeepSeek, built for robust general-purpose language understanding.
  • DeepSeek-R1 (DeepSeek-AI, 2025) is an open-source reasoning LLM from DeepSeek, trained with reinforcement learning to enhance reasoning and alignment.
  • QwQ-32b (Qwen, 2025a) is an open-source reasoning LLM by Qwen team, leveraging reinforcement learning and agent-driven tool use for complex problem solving.
  • Qwen-3-235b (Qwen, 2025b) is an open-source reasoning LLM by Qwen team, pre-trained on large corpora for high-performance language and reasoning tasks.
  • Gemini-2.5-flash (Google, 2025) is a proprietary reasoning LLM from Google’s Gemini series, optimized for ultra-long context windows and rapid inference.
  • Gemini-2.5-pro (Google, 2025) is the advanced proprietary reasoning LLM in the Gemini series, offering enhanced multimodal understanding and superior reasoning abilities.
By “reasoning” and “non-reasoning”, we distinguish whether a deliberate thinking process—often incentivized by reinforcement learning with verifiable rewards (RLVR)—is triggered before the generation of the first answer token (Li et al., 2025).
@@ A.3.3 Physical Laws
Table 4 summarizes the physical laws and their corresponding metaphysical shifts. These metaphysical shifts are achieved through mutation operations applied to the original equations, curated by human experts. The curation process follows four core guidelines:
@@
Table 4: Full physical laws and their metaphysical shifts.
@@
Fourier’s Law Thermodynamics (Thermal Conduction) $Q=k\frac{A\Delta T}{d}$ $Q=k^{\prime}\frac{A\Delta T}{d^{2}}$ $Q=k^{\prime}\frac{A+\Delta T}{d^{2}}$ $Q=k^{\prime}\frac{(A+\Delta T)A^{2}\Delta T}{d^{2}}$
$Q=k^{\prime}\frac{A\Delta T}{d^{3}}$ $Q=k^{\prime}\frac{A+\Delta T}{d^{3}}$ $Q=k^{\prime}\frac{(A+\Delta T)^{2.5}}{d^{e}}$
$Q=k^{\prime}\frac{A(\Delta T)^{2}}{d}$ $Q=k^{\prime}\frac{(A+\Delta T)^{2}}{d}$ $Q=\frac{k^{\prime}A+(\Delta T)^{2}}{d^{e}}$
@@
Modern Physics (Nuclear Physics) $N(t)=N_{0}e^{-\lambda t}$ N(t) = $N_{0}e^{-\lambda^{\prime}(t^{0.5})}$ $N(t)=N_{0}e^{-2\lambda^{\prime}+(t^{0.5})}$ $N(t)=N_{0}e^{-(2\lambda^{\prime}+2)+(t^{0.5})}$
$N(t)=N_{0}e^{-\lambda^{\prime}(t^{e})}$ $N(t)=N_{0}e^{-\lambda^{\prime}+(t^{0.5})}$ $N(t)=N_{0}e^{-\lambda^{\prime}+3+(t^{0.5})}$
$N(t)=N_{0}^{1.5}e^{-\lambda^{\prime}t}$ $N(t)=N_{0}^{1.5}e^{-\lambda^{\prime}+(t^{0.5})}$ $N(t)=\log(N_{0}^{1.5})e^{-\lambda^{\prime}+(t^{0.5})}$
@@
Wave and Acoustics (Acoustics) $v=\sqrt{\frac{\gamma RT}{M}}$ $v=\sqrt{\frac{\gamma R^{\prime}T^{2}}{M}}$ $v=\sqrt{\frac{\gamma R^{\prime}T^{2}}{M^{1.5}}}$ $v=\sqrt{\frac{e^{\gamma}R^{\prime}T^{2}}{M^{1.5}}}$
$v=\frac{\gamma R^{\prime}T}{M}$ $v=\gamma T\left(\frac{R^{\prime}}{M}\right)^{\frac{1}{3}}$ $v=\ln(\gamma)\,T\left(\frac{R^{\prime}}{M}\right)^{\frac{1}{3}}$
@@
Hooke’s Law Mechanics (Elasticity) $F=kx$ $F=2k^{\prime}x^{2}$ $F=2(k^{\prime}+x^{2})$ $F=2\left(k^{\prime}+\sin\left(x^{2}\right)\right)$
$F=2k^{\prime 2}x$ $F=2(k^{\prime 2}+x^{3})$ $F=2\left(\sin\left(k^{\prime 2}\right)+x^{3}\right)$
$F=2k^{\prime 3}x^{2}$ $F=2(k^{\prime 3}+x^{2})$ $F=2\left(\sin\left(k^{\prime 3}\right)+x^{2}\right)$

### v2

@@ 4.1 Can LLM Agents Perform Generalizable Scientific Law Discovery?
Our benchmark proves highly challenging for models with weaker reasoning abilities. All three non-thinking LLMs (GPT-4.1-mini, GPT-4.1, and DeepSeek-V3) achieve overall symbolic accuracies below 10%. While these models show limited proficiency (20-40% accuracy) even in the simplest setting, their performance degrades precipitously in all other configurations, demonstrating that success in NewtonBench is contingent on strong, generalizable reasoning.
@@
Most reasoning models achieve overall accuracies between 30-80%, with GPT-5 and Gemini-2.5-pro distinguishing themselves at 75.9% and 65.4%, respectively. While these models all demonstrate proficiency in the easiest settings (70-100% accuracy), a substantial performance gap emerges as the difficulty escalates. In the most challenging setting, GPT-5 and Gemini-2.5-pro maintain accuracies of 40.3% and 16.7%, whereas all other reasoning models fall below 6%. This stark performance gap demonstrates that the benchmark’s difficulty controls are highly effective at stress-testing and stratifying advanced reasoning abilities.
@@ 4.2 Sensitivity to Observational Noise
To investigate robustness against noisy observations, we conducted an experiment on GPT-5-mini with four levels of Gaussian noise (0.0001, 0.001, 0.01, and 0.1). As illustrated in Figure 3, performance degrades significantly even with minimal noise, with the introduction of just a 0.0001 noise level causing a 12-16% reduction in accuracy compared to the ideal, noise-free setting. As the noise level was increased from 0.0001 to 0.1, symbolic accuracy declined proportionally, while data fidelity (RMSLE) remained relatively stable. Notably, code assistance did not affect noise robustness, with the baseline and code-assisted agents degrading similarly. This demonstrates that the symbolic accuracy of LLMs is extremely fragile to noise in observational data.
@@ 4.3 Cross-Domain Performance Disparities
Performance across the physics domains, illustrated in Figure 3, reveals a substantial difference in average accuracy, with scores ranging from 18% to 54%. Bose-Einstein Distribution, the most advanced and obscure domain in our benchmark, yields the lowest average accuracy at 18.1%. Furthermore, the performance disparities are exacerbated as system complexity increases. For instance, in the simple setting, Heat Transfer yields 68% accuracy, comparable to the easiest domain, Acoustic Velocity (60%). In the complex setting, however, accuracy for Heat Transfer plummets to 3.3%, while Acoustic Velocity remains at 45.0%. This demonstrates that the intrinsic nature of a physical domain is a primary factor governing the difficulty of model discovery. Full domain analysis is provided in Appendix E.3.
@@ 4.5 The Paradox of Code Assistance: The Exploration-Exploitation Trade-Off
To investigate this hypothesis, we adapt a common analysis approach for reasoning LLMs (wang20258020rulehighentropyminority; wang2025emergenthierarchicalreasoningllms) by identifying signature tokens associated with exploration (e.g., What if, Alternatively) and exploitation (e.g., Confirm, Verify). From these, we calculate the exploration rate (the percentage of exploration tokens among all such planning tokens; see Appendix E.4 for details). Figure 6(b) shows the results for two reasoning LLMs where reasoning traces are accessible. We observe a sharp drop in the exploration rate for Gemini-2.5-flash as the code budget increases from zero to one, while the rate remains stable for Qwen-3-235b. This suggests the performance degradation in stronger models reflects over-exploitation [^3] These observations are correlational and should be interpreted as suggestive rather than causal..
@@ 5 Related Work
Symbolic regression (SR) aims to discover mathematical formulas from data, with foundational approaches being Genetic Programming (GP) that evolve populations of candidate expressions (Koza1994; augusto2000symbolic; billard2002symbolic; doi:10.1126/science.1165893). Meanwhile, more advanced approaches have been proposed, from physics-inspired, Pareto-driven strategies (udrescu2020aifeynmanphysicsinspiredmethod; udrescu2020aifeynman20paretooptimal) to deep learning-based models (petersen2021deepsymbolicregressionrecovering; kamienny2022endtoendsymbolicregressiontransformers). The latest evolution in SR leverages the reasoning capabilities of LLMs to hypothesize and refine scientific equations (shojaee2025llmsrscientificequationdiscovery; shojaee2025llmsrbenchnewbenchmarkscientific; xia2025srscientistscientificequationdiscovery; chen2025physgymbenchmarkingllmsinteractive).
@@
The agentic capabilities of LLMs are increasingly being applied to automate the scientific process (zheng2025automationautonomysurveylarge; wei2025aiscienceagenticscience), from assisting with experimental procedures in chemistry and biomedicine (gottweis2025aicoscientist; yang2025moosechemlargelanguagemodels; Luo_2025; chen2025scienceagentbenchrigorousassessmentlanguage) to automating complex research workflows in machine learning (chan2025mlebenchevaluatingmachinelearning; huang2024mlagentbenchevaluatinglanguageagents; jiang2025aideaidrivenexplorationspace; jansen2025codescientistendtoendsemiautomatedscientific). This trend culminates in the pursuit of a fully autonomous “AI Scientist” capable of managing the entire research pipeline for open-ended discovery (lu2024aiscientistfullyautomated; yamada2025aiscientistv2workshoplevelautomated).
@@
Virtual environments for LLM agents have evolved from general-purpose sandboxes for tasks like instruction following (shridhar2021alfworldaligningtextembodied), web navigation (yao2023webshopscalablerealworldweb), and compositional planning (prasad2024adaptasneededdecompositionplanning), to specialized platforms for scientific reasoning, ranging from curriculum-based tasks (wang2022scienceworldagentsmarter5th) to open-ended discovery (jansen2024discoveryworldvirtualenvironmentdeveloping).
@@ 6 Conclusion
NewtonBench is designed exclusively for benchmarking and evaluating the scientific discovery capabilities of LLM agents in a fully contamination-free, virtual environment. The counterfactual law shifts applied to physical laws ensure the tasks are novel and not directly suitable for training LLMs to discover real-world physics. Due to the interactive, model‑system nature of the tasks, we exclude traditional symbolic regression methods that do not support model systems and LLM‑based symbolic regression pipelines whose prior reliance and workflows are incompatible with our protocol; these are outside the scope of this evaluation. We caution that the benchmark is not intended as a training dataset for developing stronger physics discovery models, but rather as a controlled tool for fair and rigorous evaluation. All shifted laws were curated by expert co‑authors (Physics Olympiad backgrounds) and cross‑validated. Furthermore, while the physics domain was selected as the most representative ground for scientific law discovery, we posit that the findings of NewtonBench are generalizable to fields such as chemistry and biology. This generalization holds because the core cognitive process—isolating variables to derive mathematical relationships—is isomorphic across disciplines. Whether elucidating reaction kinetics in chemistry or modeling enzyme dynamics in quantitative biology, the discovery process demands the same rigorous interplay of hypothesis testing and symbolic abstraction. Consequently, NewtonBench serves as a robust indicator of an agent’s potential to drive discovery across the broader spectrum of the natural sciences.
@@ A.1 LLM Details
  • GPT-4.1-mini (openai_gpt41) is a lightweight proprietary non-reasoning LLM from OpenAI, designed for efficient inference.
  • GPT-4.1 (openai_gpt41) is the latest flagship proprietary non-reasoning LLM from OpenAI.
  • o4-mini (openai_o4_mini) is a proprietary reasoning LLM developed by OpenAI, optimized for lightweight inference and efficient deployment.
  • GPT-5-mini (openai_gpt_5) is a lightweight reasoning LLM from OpenAI, aiming for improved efficiency on downstream tasks.
  • GPT-5 (openai_gpt_5) is the strongest reasoning LLM by OpenAI with advanced reasoning capabilities.
  • DeepSeek-V3 (deepseekai2025deepseekv3technicalreport) is an open-source non-reasoning LLM released by DeepSeek, built for robust general-purpose language understanding.
  • DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability) is an open-source reasoning LLM from DeepSeek, trained with reinforcement learning to enhance reasoning and alignment.
  • QwQ-32b (qwq32b) is an open-source reasoning LLM by Qwen team, leveraging reinforcement learning and agent-driven tool use for complex problem solving.
  • Qwen-3-235b (yang2025qwen3technicalreport) is an open-source reasoning LLM by Qwen team, pre-trained on large corpora for high-performance language and reasoning tasks.
  • Gemini-2.5-flash (comanici2025gemini25pushingfrontier) is a proprietary reasoning LLM from Google’s Gemini series, optimized for ultra-long context windows and rapid inference.
  • Gemini-2.5-pro (comanici2025gemini25pushingfrontier) is the advanced proprietary reasoning LLM in the Gemini series, offering enhanced multimodal understanding and superior reasoning abilities.
By “reasoning” and “non-reasoning”, we distinguish whether a deliberate thinking process—often incentivized by reinforcement learning with verifiable rewards (RLVR)—is triggered before the generation of the first answer token (li202512surveyreasoning).
@@ A.2 Model Systems
System Complexity Analysis
We analyze the system complexity of all 12 physical domains based on the following two metrics:
  • Number of Assisting Equations: The number of equations $f$ applied in model $M$ that are not $f_{target}$.
  • Number of Operations: The total number of operators (e.g., ++, *, ˆ, $\sin$) that appear in all assisting laws $f$. This metric reflects the computational complexity of the model system (excluding the target law).
As illustrated in Table 3, complex systems are associated with a higher quantity of assisting equations and operations, highlighting the computational challenges inherent in the law discovery process.
Table 3: Comparison in numbers of assisting equations and operations between simple system and complex system across all 12 physical domains.
Physical Domain Simple System Complex System
Number of
Assist Equations
Number of
Operations
Number of
Assist Equations
Number of
Operations
Newton’s Law of Universal Gravitation 3 16 4 22
Coulomb’s Law 2 8 3 14
Ampère’s Force Law 4 10 5 14
Fourier’s Law 1 4 1 6
Snell’s Law 1 2 1 4
Law of Radioactive Decay 1 2 1 3
Law of Damped Harmonic Motion 1 2 2 3
Malus’s Law 1 2 1 3
Law of Sound Speed in Ideal Gas 1 2 2 7
Hooke’s Law 2 5 3 6
Bose-Einstein Distribution 1 2 2 6
Law of Heat Transfer 3 7 4 10
Average 1.75 5.17 2.42 8.17
@@ A.3.3 Physical Laws
Table 5 summarizes the physical laws and their corresponding counterfactual shifts. These counterfactual shifts are achieved through mutation operations applied to the original equations, curated by human experts. The curation process follows four core guidelines:
@@
Table 5: Full physical laws and their counterfactual shifts.
@@
Fourier’s Law Thermodynamics (Thermal Conduction) $Q=k\frac{A\Delta T}{d}$ $Q=k^{\prime}\,\dfrac{A\,\Delta T}{d^{2}}$ $Q=k^{\prime}\,\dfrac{A^{0.5}\,\Delta T}{d^{2}}$ $Q=k^{\prime}\,\dfrac{A^{0.5}\,(\Delta T)^{1.3}}{d^{2}}$
$Q=k^{\prime}\,\dfrac{A^{0.5}\,\Delta T}{d}$ $Q=k^{\prime}\,\dfrac{A^{0.5}\,(\Delta T)^{2.7}}{d}$ $Q=k^{\prime}\,\dfrac{A^{0.5}\,(\Delta T)^{2.7}}{d^{(3/7)}}$
$Q=k^{\prime}\,\dfrac{A\,(\Delta T)^{2}}{d}$ $Q=k^{\prime}\,\dfrac{(\Delta T)^{2}}{d\,A^{3.4}}$ $Q=k^{\prime}\,\dfrac{(\Delta T)^{2}}{\sqrt{d}\,A^{3.4}}$
@@
Modern Physics (Nuclear Physics) $N(t)=N_{0}e^{-\lambda t}$ $N(t)=N_{0}e^{-\lambda^{\prime}t^{1.5}}$ $N(t)=N_{0}^{1.2}e^{-\lambda^{\prime}t^{1.5}}$ $N(t)=N_{0}^{1.2}e^{-\lambda^{\prime e}t^{1.5}}$
$N(t)=N_{0}e^{-\lambda^{\prime 1.5}t}$ $N(t)=N_{0}^{1.4}e^{-\lambda^{\prime 1.5}t}$ $N(t)=N_{0}^{1.4}e^{-\lambda^{\prime 1.5}t^{e}}$
$N(t)=N_{0}e^{-(\lambda^{\prime}t)^{1.5}}$ $N(t)=N_{0}^{1.8}e^{-(\lambda^{\prime}t)^{1.5}}$ $N(t)=N_{0}^{1.8}e^{-(\lambda^{\prime}t)^{e+1.5}}$
@@
Wave and Accoustics (Acoustics) $v=\sqrt{\frac{\gamma RT}{M}}$ $v=\sqrt{\frac{\gamma R^{\prime}T^{2}}{M}}$ $v=\sqrt{\frac{\gamma R^{\prime}T^{2}}{M^{1.5}}}$ $v=\sqrt{\frac{e^{\gamma}R^{\prime}T^{2}}{M^{1.5}}}$
$v=\frac{\gamma R^{\prime}T}{M}$ $v=\frac{\gamma R^{\prime}T}{M^{\frac{1}{3}}}$ $v=\frac{\ln(\gamma)R^{\prime}T}{M^{\frac{1}{3}}}$
@@
Hooke’s Law Mechanics (Elasticity) $F=kx$ $F=2k_{1}^{\prime}x^{2}$ $F=2k_{1}^{\prime}x^{2}+k_{2}^{\prime}x$ $F=2k_{1}^{\prime}x^{2}+k_{2}^{\prime}x+k_{3}^{\prime}x^{-0.5}$
$F=2k_{1}^{\prime}x^{0.5}$ $F=2k_{1}^{\prime}x^{0.5}+k_{2}^{\prime}x^{3}$ $F=2k_{1}^{\prime}x^{0.5}+k_{2}^{\prime}x^{3}+k_{3}^{\prime}x^{-0.3}$
$F=2k_{1}^{\prime}x^{3.4}$ $F=2k_{1}^{\prime}x^{3.4}+k_{2}^{\prime}x^{0.5}$ $F=2k_{1}^{\prime}x^{3.4}+k_{2}^{\prime}x^{0.5}+k_{3}^{\prime}x^{-\frac{10}{3}}$